OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!!!! Terrence normally uses seven-eighths teaspoon of dried basil to make his own salad dressing. He decides to change the recipe to use less dried basil. He decides that the flavor is perfect when using a total of three-fourths teaspoon. Write a subtraction equation that could be used to determine the amount of basil, b, Terrence excluded. Solve for b. Describe your steps

OpenStudy (anonymous):

pla help?

OpenStudy (anonymous):

plz?

OpenStudy (anonymous):

3/4 = 7/8 - b Adding b to both sides 3/4 + b = 7/8 subtracting 3/4 from both sides b = 7/8 - 3/4 b = 1/8 1/8 basil was excluded

OpenStudy (anonymous):

it that it?

OpenStudy (anonymous):

Oui :) simple isn't it?

OpenStudy (anonymous):

yea..thank you so much..

OpenStudy (anonymous):

My pleasure Princess buddy :)

OpenStudy (anonymous):

lol :D